OpenStudy (girlgoyleh.):

WILL FAN AND MEDAL! Complete the standard form of the equation -3y = 8-4x.

OpenStudy (anonymous):

Standard form of a linear equation is Ax + By = C, so you just need to move the x-term over to the left side of the equation. Right now its -4x, so you need to add 4x to both sides. -3y = 8 - 4x +4x +4x That will cancel the x-term on the right side, and make the equation say: 4x - 3y = 8 This equation is now in standard form: Ax + By = C

OpenStudy (girlgoyleh.):

it was 4x+(-3)y=8 but thats alright